OpenStudy (anonymous):

Find y' by implicit differentiation, cos(x) + √y = 7

zepdrix (zepdrix):

So let's take the derivative of each side. Derivative of 7 with respect to x?

zepdrix (zepdrix):

\(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

sin(x\[\frac{ \sin(x) }{ \frac{ 1 }{ 2 }y ^{\frac{ -1 }{2 }} }\])

hartnn (hartnn):

thats correct, and if you need derivative only in 'x' use the original equation \(\sqrt y = 7-\cos x \)